# Fermionic Partial Transpose in the Overlap Matrix Framework for Entanglement Negativity

### v2

@@ III.2 Two numerical examples
        If we define the relation of the eigenvalues $P_{\gamma}$ and its contribution to logarithmic negativity as $f\left(P_{\gamma}\right)$, i.e. $f\left(P_{\gamma}\right)=\ln\left(1+2\sqrt{P_{\gamma}\left(1-P_{\gamma}\right)}\right)$, then the logarithmic negativity formula becomes
@@
    where the integral contour is shown in Fig. S2, and $D_{A}\left(\lambda\right)=\det\left(\tilde{M}_{A}=\lambda\mathds{1}-M_{A}\right)$ represents the characteristic polynomial of the overlap matrix, which is also a Toeplitz matrix. The Fisher-Hartwig conjecture[52, 53, 54, 6, 55, 56] facilitates the evaluation of the determinant of the Toeplitz matrix $\tilde{M}_{A}$ revealing that entanglement follows a volume law in a one-dimensional infinite chain (see Appendix B for a detailed derivation)
@@
    where $k_{F}=\frac{l}{L}\cdot 2\pi$. When $l\ll L$, the logarithmic negativity follows $\mathcal{E}\sim\frac{1}{2}\ln l+C$, where $C$ is a constant, consistent with Ref. [37, 36] for central charge $c=1$. This is also supported by the numerical results in Fig. 2(a). For higher-dimensional systems, such as a hypercube lattice $\left[0,L_{1}\right]\times\left[0,L_{2}\right]\times\cdots\times\left[0,L_{d}\right]$, the Widom conjecture[57] predicts that the entanglement scales as $\sim L^{d-1}\log L$ when region $A$ is chosen as $\left[0,z_{1}L\right]\times\left[0,z_{2}\right]\times\cdots\times\left[0,z_{d}L\right]$ with $z_{i}<1$ for $i=1,\cdots,d$. This result also aligns with previous studies[58, 54, 59].
  2. b.
    Honeycomb lattice. From Fig. 2(b), the logarithmic negativity of the honeycomb lattice follows the scaling law $\sim L\log L$, suggesting that the Widom conjecture remains valid. The scaling factor varies depending on the shape and structure of the Fermi surface. This demonstrates that both examples are consistent with the Gioev-Klich-Widom scaling law [58, 57, 60].
@@ IV Logarithmic negativity of free fermions system in tripartite case
Similar to Ref. [47], when the three overlap matrices $M_{A_{1}},\ M_{A_{2}}$ and $M_{B}$ are simultaneously diagonalizable, the logarithmic negativity
$\mathcal{E}$ $=\sum_{\gamma}\ln\left[P_{1\gamma}+P_{2\gamma}\right.$ (26)
$\hskip 36.00006pt+\sqrt{\frac{1}{2}\left(P_{B\gamma}^{2}+2P_{1\gamma}P_{2\gamma}+P_{B\gamma}\sqrt{P_{B\gamma}^{2}+4P_{1\gamma}P_{2\gamma}}\right)}$
$\hskip 36.00006pt\left.+\sqrt{\frac{1}{2}\left(P_{B\gamma}^{2}+2P_{1\gamma}P_{2\gamma}-P_{B\gamma}\sqrt{P_{B\gamma}^{2}+4P_{1\gamma}P_{2\gamma}}\right)}\right],$
where $P_{B\gamma}$ represents the probability of the $\gamma$-th particle in region $B$. It is worth noting that the complexity of this formula arises from the non-Hermitian nature of $\rho_{A}^{T_{A_{1}}^{f}}$ and simplifies to Eq. (18) when region $B$ contains no lattice sites, leading to $P_{B\gamma}=0$ for all $\gamma$.
Figure 3: Comparison of the eigenvalues of $\rho_{A}$, after reordering and removing the zero components, between the partial trace in overlap matrix approach and in Fock space. We consider an $N=12$ and $\mu/t=1$ homogeneous one-dimensional chain in two bipartite geometries: (a) an adjacent interval case where the system is evenly divided into regions $A$ and $B$; (b) a disjoint interval case where the system is evenly partitioned into $A_{1}$, $B_{1}$, $A_{2}$ and $B_{2}$ sequentially.
When three overlap matrices cannot be simultaneously diagonalized, rather than dealing with the complexity of deriving $\rho_{A}$, it is more convenient to begin the tripartite entanglement negativity calculation from Eq. (10). However, this equation should be considered rigorously, as the eigenvalues of $\rho_{A}$ have not yet been verified.
IV.1 Overlap matrix approach preserves the eigenvalues of $\rho_{A}$
In this subsection, we compare the eigenvalues of Eq. (10) with exact diagonalization results. We consider two types of tripartite geometries: adjacent intervals and disjoint intervals. As shown in Fig. 3, both geometries yield identical eigenvalues for the partial trace in different representations. Additionally, the reduced density matrix has no more than $2^{M}$ nonzero eigenvalues contributing to the entanglement. This result further demonstrates that reduced density matrices in the overlap matrix approach and Fock space are related by a unitary transformation, a property that does not always hold in the Jordan-Wigner transformation [61, 62, 63].
The key reason behind the validity of the partial trace in the overlap matrix approach lies in its “semi-local” nature, meaning that it remains associated with either region $A$ or $B$. Consequently, tracing out region $A$ or $B$ does not lead to the loss of information about the remaining region. A general proof in Appendix D and the numerical benchmark confirm that the entanglement spectrum [64, 65, 66, 67] can be reliably obtained using the overlap matrix approach.
IV.2 Logarithmic negativity for tripartite system
Previous studies [47] faced an exponential complexity issue in calculating logarithmic negativity when three overlap matrices could not be simultaneously diagonalized. The Green’s function approach provides an upper bound for the logarithmic negativity of free fermion systems in the bPT framework. Using this method, we find that although the numerical results in Ref. [47] capture the entanglement scaling law, they violate this upper bound. The upper bound of the mixed-state logarithmic negativity [38, 39, 1, 37]
@@ IV.2 Logarithmic negativity for tripartite system
where $\rho_{A}^{T_{A_{1}}^{b}}$ represents the reduced density matrix after bPT. In Fig. 4, we demonstrate that the logarithmic negativity reported in Ref. [47] exceeds the upper bound in the $M=5$ case.
Figure 4: Comparison of the logarithmic negativity in Ref. [47] with its upper bound in the $M=5$ case, highlighting the necessity of a careful approach in the partial transposition procedure.
Our clarification is as follows: while $\tilde{\rho}^{T_{B}^{f}}$ in overlap matrix approach connects to $\rho^{T_{B}^{f}}$ in Fock space via a similarity transformation for pure states, this does not hold for mixed states, which are classical mixtures of pure states. The transformations of different pure states vary, meaning that their classical mixtures disrupts the similarity transformation between $\tilde{\rho}^{T_{B}^{f}}$ and $\rho^{T_{B}^{f}}$. As a result, the findings in Ref. [47] are not totally exact.
@@ V Conclusions
In this study, we have presented a detailed analysis of the overlap matrix approach for computing entanglement in free-fermion systems, with a particular emphasis on the partial trace and fermionic partial transpose operations within the overlap matrix framework. In the bipartite case, we demonstrated that for pure states, the eigenvalues of $\rho_{A}$ remain invariant under the overlap matrix approach. This property establishes the overlap matrix approach as the first known “semi-global” mapping capable of preserving the eigenvalues of $\rho_{A}$, providing a significant advantage in entanglement computations. Moreover, we derived an analytical formula for the bipartite logarithmic negativity, offering a concise and computationally efficient method for quantifying entanglement in such systems.
@@
In summary, our work highlights both the strengths and limitations of the overlap matrix approach. For bipartite entanglement, this method proves to be a more effective and reliable strategy compared to the Green’s function approach, particularly due to its ability to preserve eigenvalues under “semi-local” mappings. However, the framework for tripartite entanglement remains incomplete, as the overlap matrix approach encounters challenges when applied to mixed states. In this context, the Green’s function approach remains the preferred choice for numerical studies due to its broader applicability.
Beyond these specific findings, our study introduces a novel perspective on entanglement research. We have demonstrated that the partial trace operation in overlap matrix approach preserves the eigenvalues of the reduced density matrix, a property that has important implications for future research. This insight provides a foundation for further refinement and application of the overlap matrix approach, potentially advancing the field of entanglement computation in free-fermion systems.
Acknowledgments— We thank Fo-Hong Wang for helpful discussion. This work was supported by the National Natural Science Foundation of China (Grants No. 12447103, No. 12274289), the National Key R&D Program of China (Grant No. 2022YFA1402702, No. 2021YFA1401400), the Innovation Program for Quantum Science and Technology (under Grant No. 2021ZD0301902), Yangyang Development Fund, and startup funds from SJTU. The computations in this paper were run on the Siyuan-1 and $\pi$ 2.0 clusters supported by the Center for High Performance Computing at Shanghai Jiao Tong University.
Appendix A Supplementary material on the bipartite overlap matrix approach
  1. a.
    The range of eigenvalues of overlap matrix. Here, we will prove that the eigenvalues of the overlap matrix lie within the range $\left[0,1\right]$. To demonstrate this, it is convenient to adopt the wave function representation. After a unitary diagonalization, the overlap matrix takes the form
@@ Appendix A Supplementary material on the bipartite overlap matrix approach
    $=\sum_{kk^{\prime}}\int_{\bm{r}\in A}\mathcal{U}_{\alpha k}\phi^{*}_{k}\left(\bm{r}\right)\phi_{k^{\prime}}\left(\bm{r}\right)\left[\mathcal{U}^{{\dagger}}\right]_{k^{\prime}\beta}\text{d}\bm{r}.$
    A new set of eigenstates $\psi_{\alpha}=\sum_{k}\left[\mathcal{U}^{{\dagger}}\right]_{k\alpha}\phi_{k}\left(\bm{r}\right)$ is constructed, preserving both orthogonality and normalization. Consequently, the integral of the squared modulus $\psi_{\alpha}$ over region $A$ cannot exceed $1$.
  2. b.
    The eigenvalues of $\tilde{\rho}_{i}^{T_{B}^{f}}$. Here we will prove that the eigenvalues of $\tilde{\rho}_{i}^{T_{B}^{f}}$ in Eq. (20) are identical to those of $\rho_{i}^{T_{B}^{f}}$ in Eq. (21). By rearranging the basis order, both matrices can be block diagonalized into the form shown in Fig. S1. The first two blocks, which contain nonzero elements (i.e. the red and orange blocks), are completely identical in both density matrices. Therefore, we only need to analyze the characteristic polynomial of the yellow block.
    Figure S1: The block diagonal structure of $\tilde{\rho}_{i}^{T_{B}^{f}}$ and $\rho_{i}^{T_{B}^{f}}$.
@@
        This demonstrates that the characteristic polynomial depends on the sequence $\left\{a_{1}b_{1},a_{2}b_{2},\cdots,a_{N-1}b_{N-1}\right\}$. Evidently, the values of $a_{i}b_{i}$ in $\rho_{i}^{T_{B}^{f}}$ are identical to those in $\tilde{\rho}_{i}^{T_{B}^{f}}$. Consequently, the characteristic polynomial and the eigenvalues remain unchanged.
Appendix B The Fisher-Hartwig conjecture
@@
In Eq. (23) the overlap matrix is a Toeplitz matrix which indicates that the matrix elements $\left[M_{A}\right]_{i,j}$ depend only on the value of $i-j$, i.e. $\left[M_{A}\right]_{i,j}=f\left(i-j\right)$ where the function $f$ depends on the specific context. The Toeplitz matrix, denoted as $T_{L}\left[\phi\right]$ in general, is generated by a function $\phi\left(\theta\right)$ if its entries are given by the Fourier coefficients of $\phi\left(\theta\right)$
$T_{L}\left[\phi\right]=\frac{1}{2\pi}\int_{0}^{2\pi}\phi(\theta)e^{-\mathrm{i}\left(i-j\right)\theta}\mathrm{d}\theta,\quad i,j=1,...,L-1$ (30)
@@
$\phi_{l}$ $\approx\int_{-\frac{1}{2}+\frac{l}{L}}^{\frac{1}{2}}\text{e}^{-\text{i}2\pi l\nu}\text{d}\nu$ (31)
$=\frac{1}{2\pi}\int_{-\pi+\frac{l}{L}\cdot 2\pi}^{\pi}\text{e}^{-\text{i}l\nu}\text{d}\nu$
$=\frac{1}{2\pi}\text{e}^{\text{i}l\pi}\int_{0}^{2\pi}\phi(\theta)\text{e}^{-\text{i}l\theta}\text{d}\theta,$
@@
$\phi\left(\theta\right)=\begin{cases}0&\text{if }\theta\in\left[0,k_{F}\right]\\ 1&\text{if }\theta\in\left[k_{F},2\pi\right]\end{cases},\ k_{F}=\frac{l}{L}\cdot 2\pi$ (32)
For the Toeplitz matrix $\tilde{M}_{A}=\lambda\mathds{1}-M_{A}$, the generating function is obtained by replacing 00 with $\lambda$ and $1$ with $\lambda-1$ respectively. Regarding the singularities of the generating function, it can be decomposed into a product of functions with known asymptotic behavior, which facilitates the application of the Fisher-Hartwig conjecture
@@
In our case, the jump discontinuity of the generating function is characterized by $R=2$, with parameters $\beta=-\beta_{1}=\beta_{2}=\frac{1}{2\pi\text{i}}\ln\ \frac{\lambda-1}{\lambda}$, and the singular points located at $\theta_{1}=k_{F}$ and $\theta_{2}=0$. The smooth part of the generating function is given by $\psi\left(\theta\right)=(\lambda-1)\left(\frac{\lambda-1}{\lambda}\right)^{-\frac{k_{F}}{2\pi}}$.
@@
$D_{L}=\left(\mathscr{F}\left[\psi\right]\right)^{L}\left(\prod_{i=1}^{R}L^{\alpha_{i}^{2}-\beta_{i}^{2}}\right)\mathscr{E}\left[\psi,\left\{\alpha_{i}\right\},\left\{\beta_{i}\right\},\left\{\theta_{i}\right\}\right]\text{ when }L\to\infty$ (35)
$\mathscr{F}[\psi]=\exp\left[\frac{1}{2\pi}\int_{0}^{2\pi}\ln\psi(\theta)\mathrm{~{}d}\theta\right]$ (36)
$\mathscr{E}[\psi,\{\alpha_{i}\},\{\beta_{i}\},\{\theta_{i}\}]$ $=\mathscr{E}[\psi]$ (37)
×$\hskip 10.00002pt\times\prod_{i=1}^{R}G(1+\alpha_{i}+\beta_{i})G(1+\alpha_{i}-\beta_{i})/G(1+2\alpha_{i})$
×$\hskip 10.00002pt\times\prod_{i=1}^{R}(\psi_{-}(\exp(\mathrm{i}\theta_{i})))^{-\alpha_{i}-\beta_{i}}(\psi_{+}(\exp(-\mathrm{i}\theta_{i})))^{-\alpha_{i}+\beta_{i}}$
×$\hskip 10.00002pt\times\prod_{1\leqslant i\neq j\leqslant R}(1-\exp(\mathrm{i}(\theta_{i}-\theta_{j})))^{-(\alpha_{i}+\beta_{i})(\alpha_{j}-\beta_{j})}$
$\mathscr{E}[\psi]=\exp\left[\sum_{k=1}^{\infty}ks_{k}s_{-k}\right]$ (38)
$G(1+z)=(2\pi)^{z/2}e^{-(z+1)z/2-\gamma_{E}z^{2}/2}\prod_{n=1}^{\infty}\{(1+z/n)^{n}e^{-z+z^{2}/(2n)}\},$ (39)
where $s_{k}$ represents the $k$th Fourier coefficient of $\ln\psi\left(\theta\right)$. In our case since $\psi\left(\theta\right)$ is a constant independent of $\theta$, we temporarily denote it as $C$, leading to $s_{k}=C\delta\left(k\right)$. Consequently, the summation simplifies as follows $\sum_{k=1}^{\infty}ks_{k}s_{-k}\rightarrow\int_{0}^{\infty}C^{2}k\delta(k)\delta(-k)\text{d}k=0$. By substituting Eq. (36), Eq. (37), Eq. (38), and Eq. (39) into Eq. (35), we arrive at the desired result.
$D_{L}(\lambda)$ $=\left[(\lambda-1)\left(\frac{\lambda-1}{\lambda}\right)^{-\frac{k_{F}}{2\pi}}\right]^{L}L^{-2\beta^{2}(\lambda)}$ (40)
$\hskip 10.00002pt\text{e}^{-(1+\gamma_{E})\beta^{2}(\lambda)}\prod_{n=1}^{\infty}\left[\left(1-\frac{\beta^{2}(\lambda)}{n^{2}}\right)^{n}\text{e}^{\frac{\beta^{2}(\lambda)}{n^{2}}}\right]$
$\hskip 10.00002pt(2-2\cos k_{F})^{-\beta^{2}(\lambda)},$

### v3

@@ III.1 One-dimensional infinite chain
  If we define the relation of the eigenvalues $P_{\gamma}$ and its contribution to logarithmic negativity as $f\left(P_{\gamma}\right)$, i.e., $f\left(P_{\gamma}\right)=\ln\left[1+2\sqrt{P_{\gamma}\left(1-P_{\gamma}\right)}\right]$, then the logarithmic negativity formula becomes
@@
where the integral contour is shown in Fig. 3, and $D_{A}\left(\lambda\right)=\det\left(\tilde{M}_{A}=\lambda\mathds{1}-M_{A}\right)$ represents the characteristic polynomial of the overlap matrix, which is also a Toeplitz matrix. The Fisher-Hartwig conjecture [63, 64, 65, 5, 66, 67] facilitates the evaluation of the determinant of the Toeplitz matrix $\tilde{M}_{A}$ revealing that entanglement follows a volume law in a one-dimensional infinite chain (see Appendix B for a detailed derivation)
@@
where $k_{F}=\frac{l}{L}2\pi$. When $l\ll L$, the logarithmic negativity follows $\mathcal{E}\sim\frac{1}{2}\ln l+C$, where $C$ is a constant, consistent with Ref. [36, 35] for central charge $c=1$. This is also supported by the numerical results in Fig. 4(a). For higher-dimensional systems, such as a hypercube lattice $\left[0,L_{1}\right]\times\left[0,L_{2}\right]\times\cdots\times\left[0,L_{d}\right]$, the Widom conjecture[68] predicts that the entanglement scales as $\sim L^{d-1}\ln L$ when region $A$ is chosen as $\left[0,z_{1}L\right]\times\left[0,z_{2}L\right]\times\cdots\times\left[0,z_{d}L\right]$ with $z_{i}<1$ for i=1,,di=1,...,d. This result also aligns with previous studies[69, 65, 70].
Figure 3: The integral contour in Eq. (42) consists of arcs $\overrightarrow{FED}$ and $\overrightarrow{CBA}$ with radii $\epsilon\to 0$. The midpoint of these arcs originate from the range of eigenvalues of the overlap matrix. The imaginary part of $\overrightarrow{AF}$($\overrightarrow{DC}$) is given by $\delta=0^{+}\left(-\delta=0^{-}\right)$.
III.2 Honeycomb lattice
From Fig. 4(b), the logarithmic negativity of the honeycomb lattice follows the scaling law $\sim L\ln L$, suggesting that the Widom conjecture remains valid. The scaling factor varies depending on the shape and structure of the Fermi surface. This demonstrates that both examples are consistent with the Gioev-Klich-Widom scaling law [69, 68, 71].
@@ IV Logarithmic negativity of free fermions system in tripartite case
Like in Ref. [49], when the three overlap matrices $M_{A_{1}},\ M_{A_{2}}$, and $M_{B}$ are simultaneously diagonalizable, their eigenvalues, denoted as $\left\{P_{A_{1}\gamma}\right\},\left\{P_{A_{2}\gamma}\right\},\left\{P_{B\gamma}\right\}$, respectively, determine the logarithmic negativity
$\mathcal{E}$ $=\sum_{\gamma}\ln\left[P_{A_{1}\gamma}+P_{A_{2}\gamma}+\sqrt{\frac{1}{2}\left(P_{B\gamma}^{2}+2P_{A_{1}\gamma}P_{A_{2}\gamma}+P_{B\gamma}\sqrt{P_{B\gamma}^{2}+4P_{A_{1}\gamma}P_{A_{2}\gamma}}\right)}\right$. (26)
$\hskip 40.00006pt\left.+\sqrt{\frac{1}{2}\left(P_{B\gamma}^{2}+2P_{A_{1}\gamma}P_{A_{2}\gamma}-P_{B\gamma}\sqrt{P_{B\gamma}^{2}+4P_{A_{1}\gamma}P_{A_{2}\gamma}}\right)}\right],$
where $P_{B\gamma}$ represents the probability of the $\gamma$-th particle in region $B$, and fulfills the identity $P_{A_{1}\gamma}+P_{A_{2}\gamma}+P_{B\gamma}=1$. A key distinction of this formula compared with that in Ref. [49] arises from the nature of the phase factor in the fermionic partial transpose. Significantly, it simplifies to Eq. (18) when region $B$ contains no lattice sites, leading to $P_{B\gamma}=0$ for all $\gamma$.
Figure 4: Logarithmic negativity $\mathcal{E}$ in different lattice systems. (a) $\mathcal{E}$ as a function of $\ln l$ in a one-dimensional infinite chain for a bipartite system. The total chain length is set to 10001000 sites. Two parameter choices are compared with the analytical formula (black line), both confirming the linear relationship between $\mathcal{E}$ and the logarithm of the subsystem size. (b) $\mathcal{E}/L$ vs $\ln L$ in a honeycomb lattice. The results indicate that logarithmic negativity follows the two-dimensional scaling law $\mathcal{E}\sim L\ln L$. The slope varies for different $\mu/t$, reflecting changes in the Fermi surface [53].
When three overlap matrices cannot be simultaneously diagonalized, previous studies [49] faced an exponential complexity issue in calculating logarithmic negativity. The Green’s function approach provides an upper bound for the logarithmic negativity of free-fermion systems in the bPT framework. Using this method, we find that although the numerical results in Ref. [49] capture the entanglement scaling law, they violate this upper bound. The upper bound of the mixed-state logarithmic negativity is [37, 38, 39, 36]
@@
where $\rho_{A}^{T_{A_{1}}^{b}}$ represents the reduced density matrix after bPT. In Fig. 5, we demonstrate that the logarithmic negativity reported in Ref. [49] exceeds the upper bound for the $M=5$ case in the one-dimensional homogeneous chain.
Figure 5: Comparison of the logarithmic negativity in Ref. [49] with its upper bound in the $M=5$ case, highlighting the necessity of a careful approach in the partial transpose procedure. We consider a one-dimensional homogeneous chain of length $L=100$ with periodic boundary conditions. The specific tripartite geometry used in our analysis is illustrated in the inset figure [53].
Our clarification is as follows: While $\tilde{\rho}^{T_{B}^{f}}$ in the overlap matrix approach connects to $\rho^{T_{B}^{f}}$ in Fock space via a similarity transformation for pure states, this does not hold for mixed states, which are classical mixtures of pure states. The transformations of different pure states vary, meaning that their classical mixtures disrupt the similarity transformation between $\tilde{\rho}^{T_{B}^{f}}$ and $\rho^{T_{B}^{f}}$. As a result, the findings in Ref. [49] are not totally exact.
@@ V Conclusions
In this study, we have presented a detailed analysis of the overlap matrix approach for computing the pure-state entanglement in free-fermion systems, with a particular emphasis on the partial trace and fermionic partial transpose operations within the overlap matrix framework. In the bipartite case, we demonstrated that for pure states, the eigenvalues of $\rho_{A}$ remain invariant under the overlap matrix approach. This property establishes the overlap matrix approach as a semiglobal mapping capable of preserving the eigenvalues of $\rho_{A}$, providing a significant advantage in entanglement computations. Moreover, we derived an analytical formula for the bipartite logarithmic negativity, offering a concise and computationally efficient method for quantifying entanglement in such systems.
Figure 6: The block diagonal structure of $\tilde{\rho}_{i}^{T_{B}^{f}}$ and $\rho_{i}^{T_{B}^{f}}$.
@@
In summary, in this work, we highlight both the strengths and limitations of the overlap matrix approach. For bipartite entanglement, this method proves to be a more effective and reliable strategy than the Green’s function approach, particularly due to its ability to preserve eigenvalues under semilocal mappings. However, the framework for tripartite entanglement remains incomplete, as the overlap matrix approach encounters challenges when applied to mixed states. In this context, the Green’s function approach remains the preferred choice for numerical studies due to its broader applicability.
Beyond these specific findings, in this study, we introduce an additional perspective on entanglement research. We have demonstrated that the partial trace operation in the overlap matrix approach preserves the eigenvalues of the reduced density matrix, a property that has important implications for future research. This insight provides a foundation for further refinement and application of the overlap matrix approach, potentially advancing the field of entanglement computation in free-fermion systems.
Acknowledgments
We thank Fo-Hong Wang for helpful discussion. This work was supported by the National Natural Science Foundation of China (Grants No. 12447103 and No. 12274289), the National Key R&D Program of China (Grants No. 2022YFA1402702 and No. 2021YFA1401400), the Innovation Program for Quantum Science and Technology (under Grant No. 2021ZD0301902), the Yangyang Development Fund, Shanghai Jiao Tong University 2030 Initiative, and startup funds from SJTU. The computations in this paper were run on the Siyuan-1 and $\pi$ 2.0 clusters supported by the Center for High Performance Computing at Shanghai Jiao Tong University.
data availability
The data that support the findings of this article are openly available [53], embargo periods may apply.
Appendix A Bipartiteoverlapmatrixapproach
A.1 The range of eigenvalues of overlap matrix.
Here, we will prove that the eigenvalues of the overlap matrix lie within the range $\left[0,1\right]$. To demonstrate this, it is convenient to adopt the wave function representation. After a unitary diagonalization, the overlap matrix takes the form
@@ Appendix A Bipartiteoverlapmatrixapproach
$=\sum_{kk^{\prime}}\int_{\mathbf{r}\in A}\mathcal{U}_{\alpha k}\phi^{*}_{k}\left(\mathbf{r}\right)\phi_{k^{\prime}}\left(\mathbf{r}\right)\left[\mathcal{U}^{{\dagger}}\right]_{k^{\prime}\beta}\text{d}\mathbf{r}.$
A new set of eigenstates $\psi_{\alpha}=\sum_{k}\left[\mathcal{U}^{{\dagger}}\right]_{k\alpha}\phi_{k}\left(\mathbf{r}\right)$ is constructed, preserving both orthogonality and normalization. Consequently, the integral of the squared modulus $\psi_{\alpha}$ over region $A$ cannot exceed $1$.
A.2 The eigenvalues of $\tilde{\rho}_{i}^{T_{B}^{f}}$.
Here we will prove that the eigenvalues of $\tilde{\rho}_{i}^{T_{B}^{f}}$ in Eq. (20) are identical to those of $\rho_{i}^{T_{B}^{f}}$ in Eq. (21). By rearranging the basis order, both matrices can be block diagonalized into the form shown in Fig. 6. The first two blocks, which contain nonzero elements (i.e., the red and orange blocks), are completely identical in both density matrices. Therefore, we only need to analyze the characteristic polynomial of the yellow block.
@@
  This demonstrates that the characteristic polynomial depends on the sequence $\left\{a_{1}b_{1},a_{2}b_{2},...,a_{N-1}b_{N-1}\right\}$. Evidently, the values of $a_{i}b_{i}$ in $\rho_{i}^{T_{B}^{f}}$ are identical to those in $\tilde{\rho}_{i}^{T_{B}^{f}}$. Consequently, the characteristic polynomial and the eigenvalues remain unchanged.
Appendix B Fisher-Hartwig conjecture
@@
In Eq. (23), the overlap matrix is a Toeplitz matrix which indicates that the matrix elements $\left[M_{A}\right]_{i,j}$ depend only on the value of iji-j, i.e., $\left[M_{A}\right]_{i,j}=f\left(i-j\right)$, where the function $f$ depends on the specific context. The Toeplitz matrix, denoted as $T_{L}\left[\phi\right]$ in general, is generated by a function $\phi\left(\theta\right)$ if its entries are given by the Fourier coefficients of $\phi\left(\theta\right)$
$T_{L}\left[\phi\right]=\frac{1}{2\pi}\int_{0}^{2\pi}\phi(\theta)\exp\left[-\mathrm{i}\left(i-j\right)\theta\right]\mathrm{d}\theta,\quad i,j=1,\ldots,L-1.$ (30)
@@
$\phi_{l}$ $\approx\int_{-\frac{1}{2}+\frac{l}{L}}^{\frac{1}{2}}\exp\left(-\text{i}2\pi l\nu\right)\text{d}\nu$ (31)
$=\frac{1}{2\pi}\int_{-\pi+\frac{l}{L}\cdot 2\pi}^{\pi}\exp\left(-\text{i}l\nu\right)\text{d}\nu$
$=\frac{1}{2\pi}\text{e}^{\text{i}l\pi}\int_{0}^{2\pi}\phi(\theta)\exp\left(-\text{i}l\theta\right)\text{d}\theta,$
@@
$\phi\left(\theta\right)=\begin{cases}0&\text{if }\theta\in\left[0,k_{F}\right]\\ 1&\text{if }\theta\in\left[k_{F},2\pi\right]\end{cases},\ k_{F}=\frac{l}{L}2\pi$ (32)
For the Toeplitz matrix $\tilde{M}_{A}=\lambda\mathds{1}-M_{A}$, the generating function is obtained by replacing 0 with $\lambda$ and $1$ with $\lambda-1$. Regarding the singularities of the generating function, it can be decomposed into a product of functions with known asymptotic behavior, which facilitates the application of the Fisher-Hartwig conjecture
@@
In our case, the jump discontinuity of the generating function is characterized by $R=2$, with parameters $\beta=-\beta_{1}=\beta_{2}=\frac{1}{2\pi\text{i}}\ln\ \frac{\lambda-1}{\lambda}$ and the singular points located at $\theta_{1}=k_{F}$ and $\theta_{2}=0$. The smooth part of the generating function is given by $\psi\left(\theta\right)=(\lambda-1)\left(\frac{\lambda-1}{\lambda}\right)^{-k_{F}/2\pi}$.
@@
$D_{L}=\left(\mathscr{F}\left[\psi\right]\right)^{L}\left(\prod_{i=1}^{R}L^{\alpha_{i}^{2}-\beta_{i}^{2}}\right)\mathscr{E}\left[\psi,\left\{\alpha_{i}\right\},\left\{\beta_{i}\right\},\left\{\theta_{i}\right\}\right]\text{ when }L\to\infty,$ (35)
$\mathscr{F}[\psi]=\exp\left[\frac{1}{2\pi}\int_{0}^{2\pi}\ln\psi(\theta)\mathrm{~d}\theta\right],$ (36)
$\mathscr{E}[\psi,\{\alpha_{i}\},\{\beta_{i}\},\{\theta_{i}\}]$ $=\mathscr{E}[\psi]\prod_{i=1}^{R}\frac{G(1+\alpha_{i}+\beta_{i})G(1+\alpha_{i}-\beta_{i})}{G(1+2\alpha_{i})}\prod_{i=1}^{R}(\psi_{-}(\exp(\mathrm{i}\theta_{i})))^{-\alpha_{i}-\beta_{i}}(\psi_{+}(\exp(-\mathrm{i}\theta_{i})))^{-\alpha_{i}+\beta_{i}}$ (37)
$\hskip 10.00002pt\times\prod_{1\leqslant i\neq j\leqslant R}(1-\exp(\mathrm{i}(\theta_{i}-\theta_{j})))^{-(\alpha_{i}+\beta_{i})(\alpha_{j}-\beta_{j})},$
$\mathscr{E}[\psi]=\exp\left[\sum_{k=1}^{\infty}ks_{k}s_{-k}\right],$ (38)
$G(1+z)=(2\pi)^{z/2}\exp\left[-\frac{(z+1)z}{2}-\frac{\gamma_{E}z^{2}}{2}\right]\prod_{n=1}^{\infty}\left\{\left(1+\frac{z}{n}\right)^{n}\exp\left(-z+\frac{z^{2}}{2n}\right)\right\},$ (39)
where $s_{k}$ represents the $k$th Fourier coefficient of $\ln\psi\left(\theta\right)$. In our case since $\psi\left(\theta\right)$ is a constant independent of $\theta$, we temporarily denote it as $C$, leading to $s_{k}=C\delta\left(k\right)$. Consequently, the summation simplifies as follows $\sum_{k=1}^{\infty}ks_{k}s_{-k}\rightarrow\int_{0}^{\infty}C^{2}k\delta(k)\delta(-k)\text{d}k=0$. By substituting Eqs. (36)-(39) into Eq. (35), we arrive at the desired result.
$D_{L}(\lambda)$ $=\left[(\lambda-1)\left(\frac{\lambda-1}{\lambda}\right)^{-k_{F}/2\pi}\right]^{L}L^{-2\beta^{2}(\lambda)}\exp\left[-(1+\gamma_{E})\beta^{2}(\lambda)\right]\prod_{n=1}^{\infty}\left\{\left[1-\frac{\beta^{2}(\lambda)}{n^{2}}\right]^{n}\exp\left[\frac{\beta^{2}(\lambda)}{n^{2}}\right]\right\}(2-2\cos k_{F})^{-\beta^{2}(\lambda)},$ (40)